Formulate the situation as a system of two linear equations in two variables. Be sure to state clearly the meaning of your x- and y-variables. Solve the system by the elimination method. Be sure to state your final answer in terms of the original question. A lawyer has found 60 investors for a limited partnership to purchase an inner-city apartment building, with each contributing either $5,000 or $10,000. If the partnership raised $435,000, then how many investors contributed $5,000 and how many contributed $10,000? $5,000 investors $10,000 investors
Answer #1

As the total number of investors are 60 => x + y = 60

Also, 5,000 * x + 10,000 * y = 435,000

=> x + 2y = 87

=> y = 87 - 60 = 27

and x = 60 - 27 = 33